OpenStudy (anonymous):

how come 10/square root 5 is the same as 2/square root 5?

jimthompson5910 (jim_thompson5910):

\[\large \frac{10}{\sqrt{5}}\neq \frac{2}{\sqrt{5}}\]

OpenStudy (anonymous):

oh wait i get it, you just rationalise it

jimthompson5910 (jim_thompson5910):

so there must be a typo somewhere

OpenStudy (anonymous):

oh sorry i meant 2squareroot 5

OpenStudy (saifoo.khan):

\[\frac{\sqrt{10}}{\sqrt{5}}=\sqrt{\frac{10}{5}}\]

jimthompson5910 (jim_thompson5910):

oh ok, yeah rationalize the denominator by multiplying both top and bottom by sqrt(5) to get your answer

OpenStudy (zarkon):

\[\frac{10}{\sqrt{5}}=\frac{10}{\sqrt{5}}\frac{\sqrt{5}}{\sqrt{5}}=\frac{10\sqrt{5}}{5}=2\sqrt{5}\]

OpenStudy (anonymous):

thank you